OpenStudy (anonymous):

find the real numbers x and y that make the equation true -3 + yi = x + 6i

hartnn (hartnn):

just compare real and imaginary parts like , if a+bi = m+ni, then a=m, b=n

hartnn (hartnn):

so, what about , -3 + yi = x + 6i ??

hartnn (hartnn):

bdw, @meow101 \(\huge \color{red}{ Welcome \: \: to \: \: Open Study \ddot \smile }\)